OpenStudy (anonymous):

Solve: 4/7 +3-(x+3) = 5/3

OpenStudy (anonymous):

we have : 7/7-x-3=5/3 Right ?

OpenStudy (anonymous):

oops I meant 4/7 + 3/(x+3) =5/3

OpenStudy (anonymous):

So it s that but easyer !

OpenStudy (anonymous):

THEN we have : -x=8/3-7/7 -x=0/6 x=-0/6

OpenStudy (anonymous):

how did you get -x?

OpenStudy (anonymous):



OpenStudy (anonymous):

Oh ok I understand now, thank you!

OpenStudy (anonymous):